OpenStudy (anonymous):

What is 2 square root of 9? Help please.

jimthompson5910 (jim_thompson5910):

What is the square root of 9?

OpenStudy (anonymous):

3.

jimthompson5910 (jim_thompson5910):

so \[\Large 2\sqrt{9} = 2*3 = 6\]

OpenStudy (anonymous):

Alright! So I was right. Thanks!

jimthompson5910 (jim_thompson5910):

sure thing

jimthompson5910 (jim_thompson5910):

No, it's just 3